Question 4

Orville Company’s standard and actual costs per unit are provided below for the most recent period. During this time period 400 units were actually produced.

Standard

Actual

Materials:

     Standard: 2 metres at $1.50 per m.

$3.00

     Actual: 2.1 metres at $1.60 per m.

$3.36

Direct labour:

     Standard: 1.5 hrs. at $6.00 per hr.

9.00

     Actual: 1.4 hrs. at $6.50 per hr.

9.10

Variable overhead:

     Standard: 1.5 hrs. at $3.40 per hr.

5.10

     Actual: 1.4 hrs. at $3.10 per hr.

4.34

Total unit cost

$17.19

$16.80

For simplicity, assume there was no inventory of materials at the beginning or end of the period.

Required:

Given the information above, compute the following variances. Also indicate if the variances are favorable or unfavorable.

  1. Materials price variance
  2. Materials quantity variance
  3. Direct labour rate variance
  4. Direct labour efficiency variance
  5. Variable overhead spending variance
  6. Variable overhead efficiency variance
Answer #1

Material Price Variance = (Actual Price - Standard Price) x Actual Quantity
= ($1.60 - $1.50) x 840 = $84 (U)
Actual Quantity = 400 x 2.1 = 840

Material Quantity Variance = (Actual Quantity - Standard Quantity) x Standard Price
= (840 - 800) x $1.50 = $60 (U)
Standard Quantity = 400 x 2 = 800

Direct Labor rate variance = (Actual rate - Standard rate) x actual hours
= ($6.50 - $6) x 560 = $380 (U)
Actual hours = 400 x 1.4 = 560 hours

Direct labor efficiency variance = (Actual hours - standard hours) x Standard rate
= (600 - 560) x $6 = $240 (F)
Standard hours = 1.5 x 400 = 600

Variable overhead spending variance = (Actual rate - Standard rate) x actual hours
= ($3.10 - $3.40) x 560 = $168 (F)
Actual hours = 400 x 1.4 = 560 hours

Direct labor efficiency variance = (Actual hours - standard hours) x Standard rate
= (600 - 560) x $3.40 = $136 (F)